OpenStudy (anonymous):

Calculus and Vectors: (Step by Step please - I really don't understand this one) Let u = 4i-2j+ 3k and v = [−1, 0, 3]. Determine the following. a. 3u - 2v b. | v| c. directional angles of v d. unit vector of v .

OpenStudy (anonymous):

u =<4,-2,3> and v=<-1,0,3> so 3u =<12,-6,9> and 2v= <-2,0,6> hence 3u-2v = 10,-6,3>

OpenStudy (anonymous):

|v|=sqrt( 1+0+9)=sqrt(10)

OpenStudy (anonymous):

direction angles of v are cos^-1(-1/sqrt(10)) , 90 degree, cos^-1(-3/sqrt(10)) plz neglect any typo error

OpenStudy (anonymous):

unit vector of v = <-1/sqrt(10),0,3/sqrt(10)>

OpenStudy (anonymous):

wouldn't the answer be (14,-6,3) for the first one?

OpenStudy (dumbcow):

yes its (14,-6,3) , they just had a typo

OpenStudy (anonymous):

alright, than kyou